OpenStudy (anonymous):

Ur, another integration-by-parts problem! I have g(a) = the integral from -a to a of e^((-x^2)/9) (x^4)dx. I know two things: I need to integrate twice, and the answer will model the erf(q) equation.

OpenStudy (anonymous):

But what I don't get is how to split up the variables to do my intergration-by-parts. Is it u= x^4, dv= e^((-x^2)/2), or vice versa?

OpenStudy (anonymous):

either way, I have been struggling with this and not getting a satisfactory answer

OpenStudy (anonymous):

Definitely make u equal to the x^4 ! Use the acronym LIATE to figure out what should be u, so that way you don't always guess it out. L - logarithmic I - inverse trig A - algebraic T - trig E - exponential So since algebraic comes before exponential, use the x^4 as your u.

OpenStudy (turingtest):

I think we can make use of the same trick we did last time since this function is even

OpenStudy (turingtest):

yes we can, we just have to do it twice

OpenStudy (anonymous):

Hi guys, just left for a quick sec

OpenStudy (anonymous):

So, I was looking at that: can I split the x^4 up? I was trying, and making two intergrations at the moment

OpenStudy (anonymous):

but I get\[\int\limits_{-a}^{a}w^2 *e^-w/9\]?

OpenStudy (anonymous):

I'm guessing I can do u = w^2, du = 2w; dv = e^(-w/9), v = e^(-w/9) at this point?

OpenStudy (anonymous):

that's what I'm trying right now...

OpenStudy (turingtest):

\[\Large\int_{-a}^ax^4e^{-\frac{x^2}9}dx=\int_0^ax^3(2xe^{-\frac{x^2}9})dx\]\[u=x^3\implies du=3x^2dx\]\[dv=2xe^{-\frac{x^2}9}dx\implies v=-9e^{-\frac{x^2}9}\]\[\int_0^ax^3(2xe^{-\frac{x^2}9})dx=-9x^3e^{-\frac{x^2}9}+27\int_0^ax^2e^{-\frac{x^2}9}dx\]\[=-9x^3e^{-\frac{x^2}9}+27\int_0^ax(xe^{-\frac{x^2}9})dx\]\[u=x\implies du=dx\]\[dv=xe^{-\frac{x^2}9}dx\implies v=-\frac92e^{-\frac{x^2}9}\]so the integral is now\[-9x^3e^{-\frac{x^2}9}+27(-\frac92xe^{-\frac{x^2}9}+\frac92\int e^{-\frac{x^2}9}dx)\]crap I got stuck...

OpenStudy (anonymous):

wait, that's fine I think: I need to integrate twice, and then write it in terms of the erf(q) function I mentioned above. I have similar work like yours, up to a point...

OpenStudy (turingtest):

oh you can use the error function ?! I missed that the last integral is the error function

OpenStudy (turingtest):

so then my answer is right :)

OpenStudy (anonymous):

Wait, so this is how I reasoned the first substution: did you do something like w = x^2, dw = 2xdx? and then so it would be the integeral 2xwe^-wdx before you did intergration by parts?

OpenStudy (turingtest):

no I didn't think like that I used the even function thing to get it into the form I have on the first line then did u=x^3, dv=2xe^(-x^2/9)

OpenStudy (anonymous):

hmm..

OpenStudy (anonymous):

I guess I'm having trouble intergrating e^((-x^2)/9) once I try to do the first intergration by parts

OpenStudy (turingtest):

\[\Large\int_{-a}^ax^4e^{-\frac{x^2}9}dx=\int_0^ax^3(2xe^{-\frac{x^2}9})dx\]\[u=x^3\implies du=3x^2dx\]\[dv=2xe^{-\frac{x^2}9}dx\implies v=-9e^{-\frac{x^2}9}\]\[\int_0^ax^3(2xe^{-\frac{x^2}9})dx=-9a^3e^{-\frac{a^2}9}+27\int_0^ax^2e^{-\frac{x^2}9}dx\]\[=-9a^3e^{-\frac{a^2}9}+27\int_0^ax(xe^{-\frac{x^2}9})dx\]\[u=x\implies du=dx\]\[dv=xe^{-\frac{x^2}9}dx\implies v=-\frac92e^{-\frac{x^2}9}\]so the integral is now\[-9a^3e^{-\frac{a^2}9}+27(-\frac92ae^{-\frac{a^2}9}+\frac92\int_0^a e^{-\frac{x^2}9}dx)\]\[-9a^3e^{-\frac{a^2}9}+27\left(-\frac92ae^{-\frac{a^2}9}+\frac{27}4\sqrt\pi\text{erf}(\frac a3)\right)\]I'll let you simplify that ;)

OpenStudy (turingtest):

okay your problem is that you \(cannot\) integrate\[\int e^{-\frac{x^2}9}dx\]in terms of traditional functions, the answer to the integral is\[\int e^{-\frac{x^2}9}dx=\frac32\sqrt\pi\text{erf}(\frac x3)\]so that is where you are supposed to be stuck :P

OpenStudy (anonymous):

right; so I am having issues with doing intergration by parts while being stuck there.

OpenStudy (anonymous):

It seems that you don't try to intergrate it at all, you just kind of go with it

OpenStudy (turingtest):

you can prove that integral is what it is if you know a bit about the gamma function and laplace transforms

OpenStudy (anonymous):

But you still take the intergal of dv...how did you do the integral of 2xe^((-x^2)/9)?

OpenStudy (turingtest):

let \[u=e^{-\frac{x^2}9}\implies du=-\frac29xe^{-\frac{x^2}9}dx\]integral becomes\[-9\int du\]

OpenStudy (turingtest):

gotta do those simple u-subs in your head as you move through calculus

OpenStudy (anonymous):

yea, I get that so far

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

\[\int\limits_{}^{} e^(x^2/9) dx\] = ? (2x/9)*e^((-x^2)/9)?

OpenStudy (anonymous):

I don't think it's as easy as that---or I'm doing something wrong

OpenStudy (anonymous):

wait, you don't intergrate/derive the x^2 anymore, only the 1/9 in your work

OpenStudy (turingtest):

I am confused what you mean ...btw to keep the exponents where they are supposed to be enclose them in {} not in ()

OpenStudy (turingtest):

as I said before you cannot integrate e^(-x^2/9)dx, but you can integrate 2xe^(-x^2/9)dx the 2x outside allows us to do a simple substitution

OpenStudy (anonymous):

oh!

OpenStudy (turingtest):

the trick is realizing how similar 2xe^(-x^2/9) is to the derivative of e^(-x^2/9)

OpenStudy (turingtest):

I think it clicked, yes?

OpenStudy (turingtest):

I'll brb...

OpenStudy (anonymous):

I got it! Thanks so much! It finally clicked!

OpenStudy (turingtest):

I'm glad, I think I will celebrate this rather difficult integration by parts problem with a beer see ya!

OpenStudy (anonymous):

see yah!

OpenStudy (anonymous):

The main trick is to know when to separate the powers of x to fit your purposes, I guess. This will probably appear on an exam.